OpenStudy (anonymous):

The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle? in.

OpenStudy (calculusxy):



OpenStudy (calculusxy):

\[2(8 + w) + 2(w) = 52\] solve for w

OpenStudy (anonymous):

9

OpenStudy (calculusxy):

good!

OpenStudy (anonymous):

YAY!!!!!!!!!!!!

OpenStudy (calculusxy):

Plug it into the value for \(L = 8 + w\)

OpenStudy (anonymous):

2(8+w)+2(w)=52 (16+2w)+(2w)=52 16+4w=52 4w=36 w=4

OpenStudy (anonymous):

L=8+4 L=12

OpenStudy (calculusxy):

@MJgirL1958 Sorry you're wrong ...

OpenStudy (anonymous):

So the length is 12

OpenStudy (anonymous):

ok im always wrong so ;)

OpenStudy (anonymous):

write down what you solved

OpenStudy (calculusxy):

You would get w = 4 which is what @Abdullahbasra007 did. And then you would have to do \( L = 8 + 4\) to get \(\color{green}{12}\)

OpenStudy (anonymous):

OH ok i see it now thank you both

OpenStudy (anonymous):

Anytime

OpenStudy (calculusxy):

\(\huge \scr WELCOME !\)

OpenStudy (anonymous):

Thanks for the medal

OpenStudy (anonymous):

welcome ;)